OpenStudy (anonymous):

use the table below to find (FoG)(1) x 0 1 2 3 4 5 6 7 f(x) 5 7 9 11 13 15 17 19 g(x) 3 6 9 12 15 18 21 24

OpenStudy (anonymous):

(FoG)(1) = f(g(1))

OpenStudy (anonymous):

the table is supposed to look like this

OpenStudy (anonymous):

answer is 17.. (fog)(1)=f(g(1))+f(6)=17

OpenStudy (anonymous):

x 0 1 2 3 4 5 6 7 f(x) 5 7 9 11 13 15 17 19 g(x) 3 6 9 12 15 18 21 24

OpenStudy (anonymous):

I was told we're not allowed to give answers anymore KBOY

OpenStudy (anonymous):

Like, you're supposed to lead them to the answer now....

OpenStudy (anonymous):

f(g(1))= f(6) not + f(6)..typo

OpenStudy (anonymous):

why would that be that case? thanks! for doing it

OpenStudy (anonymous):

At least give me credit for giving you the formula :/